OpenStudy (anonymous):

Can anyone please help me by solving: 3sinx - cosx = 3

OpenStudy (michele_laino):

please I can rewrite your equation, as follows, namely I substitute sin x with: \[\sqrt{1-(\cos x)^{2}}\], so: \[3\sqrt{1-(\cos x)^{2}}=3+\cos x\] is it clear, please?

OpenStudy (michele_laino):

since 3 sin x=3+cos x